# Real-time 3D-aware Portrait Editing from a Single Image

### v1

@@ 3.2. Distilling Priors into a Lightweight Module
Figure 3. Disentanglement in Live3D features. We separately disable the features from $\mathrm{E}_{high}(\cdot)$ and $\mathrm{E}_{low}(\cdot)$ to infer the reconstructed image. Without $\mathrm{E}_{high}(\cdot)$, the output retains the coarse structure but loses the appearance. Conversely, when $\mathrm{E}_{low}(\cdot)$ is deactivated, the reconstructed portraits preserve the texture (such as the blue and purple reflection on the glasses) but fail to capture the geometry.
Injecting Prompts $\mathbf{P}$ as Condition. In contrast to input images, prompts typically offer more high-frequency and texture information to guide and control the editing process. Accordingly, it is appropriate to incorporate the prompts into the high-level branch $\mathrm{E}_{high}(\cdot)$ in Live3D to extract the high-frequency features. To incorporate prompts, we utilize the prompt encoder $\mathrm{E}_{p}(\cdot)$ to create prompt embeddings that are fused into our model through cross attention, which is similar to the strategy in Stable Diffusion (Rombach et al., 2022). Specifically, we add a cross-attention layer after $\mathrm{E}_{high}(\cdot)$ to obtain the feature updated with the prompt embeddings:
(2) $\mathbf{F}_{a}$ $=\mathrm{crossatt}(\mathrm{E}_{high}(\mathbf{I}),\mathrm{E}_{p}(\mathbf{P})),$
where the encoder $\mathbf{E}_{p}(\cdot)$ is a transformer (MAE (He et al., 2022) for image and CLIP (Radford et al., 2021) for text). The $\mathrm{E}_{high}(\mathbf{I})$ serves as the query and $\mathrm{E}_{p}(\mathbf{P})$ is employed as key and value. The updated feature $\mathbf{F}_{a}$ and the geometry feature $\mathbf{F}_{g}$ are then feed into the decoder $\mathrm{E}_{t}(\cdot)$ to infer the triplane $\mathbf{T}_{\mathbf{p}}$:
(3) $\mathbf{T}_{\mathbf{p}}=\mathrm{E}_{t}(\mathbf{F}_{a},\mathbf{F}_{g}).$
Distilling Diffusion and 3D GAN Prior. Unlike previous methods that require large amounts of 3D data, our model operates with only 2D paired data generated by the 2D editing model (Brooks et al., 2023) during training. For visual prompts, we perform captioning (Li et al., 2022) on $\mathbf{P}$ and then use the captions as the instruction to generate an edited image $\mathbf{I}_{gt}$ with 2D diffusion model (Brooks et al., 2023) as the pseudo label. For text prompts, we follow a similar process but skip the captioning step. With the predicted triplane $\mathbf{T}_{\mathbf{p}}$, we can render the image using the pretrained EG3D model and then calculate the reconstruction loss as follows:
(4) $\mathcal{L}_{2d}=\ell_{\mathbf{I}}(\mathbf{I}_{gt},\mathrm{R}(\mathbf{T_{\mathbf{p}}},\mathbf{c})),$
where $\mathrm{R}(\cdot)$ is the rendering module of EG3D, $\mathbf{c}$ is the camera pose of input portrait, and $\ell_{1}(\cdot)$ is an image reconstruction loss penalizing the difference between the ground truth $\mathbf{I}_{gt}$ and the rendering $\mathbf{I}_{p}=\mathrm{R}(\mathbf{T_{\mathbf{p}}},\mathbf{c}))$. It’s important to note that $\mathcal{L}_{2d}$, used to reconstruct the $\mathbf{I}_{gt}$, essentially distill the knowledge of editing from the diffusion model.
@@
(5) $\mathbf{T}_{gt},\mathcal{D}_{gt}=\mathrm{G}(\mathbf{I}_{gt},\mathcal{C}),$
where $\mathrm{G}(\cdot)$ represents the inference process of Live3D, where $\mathcal{C}=\{\mathbf{c}_{1},..,\mathbf{c}_{n}\}$ is the camera set, and $n$ denotes the number of cameras. We also render the multi-view depths $\mathcal{D}_{p}$ and images $\mathcal{I}_{p}$ of the edited image $\mathrm{I}_{p}$ from the triplane $\mathbf{T}_{p}$ and then define the objective:
(6) $\mathcal{L}_{3d}=\ell_{\mathcal{I}}(\mathcal{I}_{p},\mathcal{I}_{gt})+\ell_{\mathbf{T}}(\mathbf{T}_{p},\mathbf{T}_{gt})+\ell_{\mathcal{D}}(\mathcal{D}_{p},\mathcal{D}_{gt}),$
@@ 3.3. Training and Inference
Training. During the training phase, we sample a triplet consisting of the input portrait $\mathbf{I}$ along with its camera pose $\mathbf{c}$, prompts $\mathbf{P}$, and the pseudo-label image $\mathbf{I}_{gt}$ generated by the 2D diffusion model. We leverage the Live3D model as a pretrained model and add an additional prompts encoder $\mathrm{E}_{p}$ to extract prompts embeddings. The overall learning objective can be described as follows:
(7) $L=\lambda_{1}L_{2D}+\lambda_{2}L_{3D},$
@@
Inference. For inference, users can provide a single portrait image and choose prompts used during our training. Our model is able to generate the edited 3D NeRF along with photorealistic view synthesis.
Customized Prompts Adaptation. To accommodate customized styles provided by users, we propose a method to adapt our pretrained encoder to novel prompts. We increase the tuning efficiency by optimizing only $\mathrm{E}_{p}(\cdot)$ and the normalization layers in $\mathrm{E}_{t}(\cdot)$ with the same learning objective Eq. 7. This method allows us to limit the training data to only 10 image pairs and the learning time to 5 minutes on a single GPU.
4. Experiments
Figure 4. Qualitative comparisons. We compare the results of several baselines with image prompts and text prompts. In each case, we include the edited portraits as well as their novel view renderings. Our method generates high-quality edited portraits with better 3D quality and alignment with the referenced prompts.
Figure 5. Novel prompt adaptation. We show the intermediate testing results at 10s, 1min, 2min and 5min during adaptation with 10 paired training images.
4.1. Experimental Setup
Training Settings. For real faces, we adopt the FFHQ dataset (Karras et al., 2019) at $512\times 512$ resolution with camera parameters aligned by EG3D (Chan et al., 2022). In order to obtain the stylized images as pseudo labels and visual prompts, we leverage InstructPix2Pix (Brooks et al., 2023) to edit the real faces. Specifically, given an image prompt, we first employ BLIP (Li et al., 2022) for captioning and use the obtained text as the instruction to synthesize an edited image with InstructPix2Pix. We totally conduct experiments on 20 styles, and for each style, we synthesize 1000 images, resulting in 20,000 image pairs for model training. For each style, we use 8 textual prompts for training and 5 textual prompts for testing. We adopt a learning rate of 5e-5 and optimize the model for 60k iterations with a batch size of 32. The entire training procedure is completed over a period of 40 hours utilizing 8 NVIDIA A100 GPUs.
Baselines. We conduct comparisons against two categories of methods that achieve analogous outcomes: 1) 3D construction coupled with 3D editing; 2) 2D editing followed by 3D reconstruction. For the first category, we implement two baseline approaches, wherein we leverage PTI (Roich et al., 2022) or Live3D (Trevithick et al., 2023) for reconstruction and augment both of them with CLIP loss (Radford et al., 2021) for editing. In the second category, we employ an instructive pixel-to-pixel editing method (Brooks et al., 2023) prior to executing 3D reconstruction via Live3D.
Evaluation Criteria. We evaluate all metrics on 100 pairs of images processed from FFHQ. We conduct a comprehensive evaluation of the editing performance in the following four aspects: identity preservation, reference alignment, 3D consistency, and inference speed. 1) Identity preservation ($\mathbf{ID_{t}}$) aims to measure the preservation of the original identity by calculating the cosine similarity between the identity feature of the input image $\mathbf{I}$ and that of the edited image $\mathbf{I}_{p}$. We use ArcFace model (Deng et al., 2019a) to extract identity features. 2) Reference alignment ($\mathbf{CLIP_{r}}$) targets at assessing the alignment of the output editing styles to the desired input prompt by computing the cosine similarity in the CLIP (Radford et al., 2021) feature space. 3) 3D consistency (3D) on the edited outputs is measured following the evaluation protocols established by EG3D (Chan et al., 2022). This involves calculating the identity similarity across multiple views. 4) Inference speed (Time) is measured on a single NVIDIA A6000 GPU with an average of 100 samples.
Table 1. Quantitative comparisons. We compare several baselines on the 100 images of FFHQ dataset. It’s important to note that we exclude $CLIP_{r}$ for PTI+CLIP and Live3D+CLIP since these models utilize CLIP for optimization. Our model excels in 3D quality and achieves a remarkable 250x improvement (compared with IP2P+Live3D) in inference speed, achieving real-time performance.
[…]
PTI+CLIP 0.11 - 0.73 360s
Live3D+CLIP 0.63 - 0.72 30s
IP2P+Live3D 0.52 0.59 0.75 10s
@@ 4.1. Experimental Setup
4.2. Efficient 3D-aware Portrait Editing
We make in-depth analysis of our efficient portrait editing system both quantitatively and qualitatively. The results are included in Tab. 1 and Fig. 4. For reference alignment, we do not report $CLIP_{r}$ for methods that leverage CLIP for optimization since it is evaluated with CLIP as well. The standout feature of our system is its efficiency. Our method achieves an inference speed of merely 40ms, which improves over 100 times compared to the fastest existing baselines, which require around 10 seconds. Because of the efficient knowledge distillation, our approach is also good in preserving the identity, adhere more closely to the reference prompts and achieve the best 3D consistency. Although the Live3D + CLIP achieve the best $ID_{t}$ score, the edited results are basically unchanged compared to the input. We suspect that Live3D does not have a latent space and loses the editing priors, making it challenging for the results of CLIP optimization to align well with prompts. Compared with PTI + CLIP methods, our model can enable precise alignment with input and prompts, while the baseline always generates low-fidelity textures, and the resulting geometry has many artifacts. In contrast to the 2D editing and subsequent 3D reconstruction pipeline (IP2P + Live3D), our system produces textures consistent with the prompt and aligns better with the input structure. For example, as the second sample in Fig. 4 shows, our model can consistently retain the hair structure, while IP2P + Live3D cannot handle this. The third sample showcases that IP2P + Live3D introduces texture artifacts on the face and cannot inherit the style of prompts very well.
@@ 4.2. Efficient 3D-aware Portrait Editing
Figure 6. Qualitative comparison for ablations on (a) the distillation loss of 3D GAN ($\mathcal{L}_{3D}$) and (b) diffusion models ($\mathcal{L}_{2D}$).
4.3. Adaptation to Customized Editing
@@
With the trained editing network, our system already supports a variety of 3D-aware editing styles. To expand the selection and better conform to users’ preferences, we offer an efficient method for fast adaptation to new prompts. Users can personalize the editing network by providing a modest set of 10 reference editing pairs. These pairs can be either handpicked from artist-created examples or produced using text-guided image editing models. The adaptation process itself is remarkably fast, requiring only about 5 minutes to accomplish the learning of the novel prompts. In Fig. 5, we present the intermediate testing results in the adaptation process. Our model can quickly master the novel-prompt knowledge in about 2 minutes. With further training (e.g., 5 minutes), the edited results become more stylized, demonstrating a trade-off between the authenticity and the stylization. Upon completion of the adaptation, the system allows users to edit testing inputs in these newly learned styles with a minimal inference time of 0.04s. This rapid performance indicates that the system is well-optimized for real-time applications, providing a seamless and efficient user experience.
Figure 7. Video editing results with 3DPE. We use text prompts ”Bronze statue” to edit the input video. Our method can accurately reconstruct challenging facial expressions and achieve high-quality novel view renderings.
4.4. Ablation Studies
$\mathcal{L}_{2D}$ for Distillation of Diffusion Models. We also study the effect of $\mathcal{L}_{2D}$. From Fig. 6, we observe that $\mathcal{L}_{2D}$ is critical for our model, especially in preserving detailed texture. Without it, the edited image tends to lose appearance information from prompts and exhibits structural artifacts around the nose.
Ablation on Novel Prompt Adaptations. To investigate the ability of our method adapted to novel prompts, we conduct an ablation study on the number of novel-prompt data pairs required for training. We choose the golden statue as the novel prompt for demonstration and perform adaptation with 2, 5, 10, 20, or 50 pairs of novel-prompt data on a single A6000 GPU. We experimentally find that our model can converge in about 5 minutes, and therefore, we report the perceptual distance between the model output and the provided ground truth at 0.1min, 1min, 2min, and 5min for comparison. As shown in Tab. 2, our model can adapt to the new style in around 2 minutes. As the number of provided data pairs increases, the model can adapt faster at the beginning.
@@ 4.4. Ablation Studies
Table 2. Ablation study on the number of data pairs used for novel prompt adaptation. We report LPIPS score after 0.1min, 1min, 2min, and 5min fine-tuning for evaluation.
@@
4.5. Applications

### v2

@@ 3.2 Distilling Priors into a Lightweight Module
Injecting Prompts $\mathbf{P}$ as Condition. In contrast to input images, prompts typically offer more high-frequency and texture information to guide and control the editing process. Accordingly, it is appropriate to incorporate the prompts into the high-level branch $\mathrm{E}_{high}(\cdot)$ in Live3D to extract the high-frequency features. To incorporate prompts, we utilize the prompt encoder $\mathrm{E}_{p}(\cdot)$ to create prompt embeddings that are fused into our model through cross attention, which is similar to the strategy in Stable Diffusion [55]. Specifically, we add a cross-attention layer after $\mathrm{E}_{high}(\cdot)$ to obtain the feature updated with the prompt embeddings:
$\mathbf{F}_{a}$ $=\mathrm{CrossAttention}(\mathrm{E}_{high}(\mathbf{I}),\mathrm{E}_{p}(\mathbf{P})),$ (2)
where the encoder $\mathbf{E}_{p}(\cdot)$ is a transformer (MAE [29] for image and CLIP [53] for text). The $\mathrm{E}_{high}(\mathbf{I})$ serves as the query and $\mathrm{E}_{p}(\mathbf{P})$ is employed as key and value. The updated feature $\mathbf{F}_{a}$ and the geometry feature $\mathbf{F}_{g}$ are then feed into the decoder $\mathrm{E}_{t}(\cdot)$ to infer the triplane $\mathbf{T}_{\mathbf{p}}$:
$\mathbf{T}_{\mathbf{p}}=\mathrm{E}_{t}(\mathbf{F}_{a},\mathbf{F}_{g}).$ (3)
Distilling Diffusion and 3D GAN Prior. Unlike previous methods that require large amounts of 3D data, our model operates with only 2D paired data generated by the 2D editing model [7]. For both visual and textual prompts, we generate edited images $\mathbf{I}_{gt}$ with the diffusion-based image editing model [7] as the pseudo labels. The details of data generation can be found in Sec. 4.1.
With the predicted triplane $\mathbf{T}_{\mathbf{p}}$, we can render the image using the pretrained EG3D model and then calculate the reconstruction loss as follows:
$\mathcal{L}_{2d}=\ell_{\mathbf{I}}(\mathbf{I}_{gt},\mathrm{R}(\mathbf{T_{\mathbf{p}}},\mathbf{c})),$ (4)
where $\mathrm{R}(\cdot)$ is the rendering module of EG3D, $\mathbf{c}$ is the camera pose of input portrait, and $\ell_{1}(\cdot)$ is an image reconstruction loss penalizing the difference between the ground truth $\mathbf{I}_{gt}$ and the rendering $\mathbf{I}_{p}=\mathrm{R}(\mathbf{T_{\mathbf{p}}},\mathbf{c}))$. It’s important to note that $\mathcal{L}_{2d}$, used to reconstruct the $\mathbf{I}_{gt}$, essentially distills the knowledge of editing from the diffusion model.
@@
$\mathbf{T}_{gt},\mathcal{D}_{gt}=\mathrm{G}(\mathbf{I}_{gt},\mathcal{C}),$ (5)
where $\mathrm{G}(\cdot)$ represents the inference process of Live3D, where $\mathcal{C}=\{\mathbf{c}_{1},..,\mathbf{c}_{n}\}$ is the camera set, and $n$ denotes the number of cameras. We also render the multi-view images $\mathcal{I}_{p}$ and depths $\mathcal{D}_{p}$ of the edited image $\mathrm{I}_{p}$ from the triplane $\mathbf{T}_{p}$ and then define the objective:
$\mathcal{L}_{3d}=\ell_{\mathcal{I}}(\mathcal{I}_{p},\mathcal{I}_{gt})+\ell_{\mathbf{T}}(\mathbf{T}_{p},\mathbf{T}_{gt})+\ell_{\mathcal{D}}(\mathcal{D}_{p},\mathcal{D}_{gt}),$ (6)
@@ 3.3 Training and Inference
Training. During the training phase, we sample a triplet consisting of the input portrait $\mathbf{I}$ along with its camera pose $\mathbf{c}$, prompts $\mathbf{P}$, and the pseudo-label image $\mathbf{I}_{gt}$ generated by the 2D diffusion model. We leverage the Live3D model as a pretrained model and add an additional prompts encoder $\mathrm{E}_{p}$ to extract prompt embeddings. The overall learning objective can be described as follows:
$L=\lambda_{1}L_{2D}+\lambda_{2}L_{3D},$ (7)
where $\lambda_{1}$ and $\lambda_{2}$ are loss weights. In our setting, $\lambda_{1}$ and $\lambda_{2}$ are both set to 1.0. For the reconstruction loss, $\ell_{\mathcal{I}}(\cdot)$, $\ell_{\mathrm{I}}(\cdot)$ are combinations of L2 loss and LPIPS loss [82], with loss weights being 1 and 2, respectively. $\ell_{\mathcal{D}}(\cdot)$ and $\ell_{\mathrm{T}}(\cdot)$ are L1 loss. Notably, based on the study on Live3D feature representation, during training only $\mathrm{E}_{p}(\cdot)$, $\mathrm{E}_{t}(\cdot)$ are set to be learnable, while other modules are frozen. It allows our model to leverage Live3D knowledge as much as possible and converge at a very fast speed.
Inference. For inference, users can provide a single 2D portrait image and choose an image prompt or text prompt. Our model is able to generate the edited 3D NeRF along with photorealistic view synthesis.
Customized Prompts Adaptation. To accommodate customized styles provided by users, we propose a method to adapt our pretrained encoder to user-defined styles. We increase the tuning efficiency by optimizing only $\mathrm{E}_{p}(\cdot)$ and the normalization layers in $\mathrm{E}_{t}(\cdot)$ with the same learning objective Eq. 7. This method allows us to limit the training data to only 10 image pairs and the learning time to 5 minutes on a single GPU.
4 Experiments
4.1 Experimental Setup
Training Settings. For real faces, we adopt the FFHQ dataset [35] at $512\times 512$ resolution with camera parameters aligned by EG3D [13]. In order to obtain the stylized images as pseudo labels, and the corresponding textual and visual prompts, we leverage the text-guided editing model InstructPix2Pix (IP2P) [7] to edit the real faces. Specifically, we manually generate and select textual prompts, and send them to IP2P together with the real images to obtain the edited 2D images. The text prompts of our training and testing set are obtained by augmenting the aforementioned textual prompts with the large language model GPT [8] following IP2P [7]. While the image prompts of our training and testing set are picked from the edited 2D images. The dataset details and samples are provided in the appendix for further investigation. We conduct experiments on 20 styles, and for each style, we synthesize 1000 images, resulting in 20,000 image pairs with text prompts for model training. For each style, we use 8 textual prompts for training and 5 textual prompts for testing. We adopt a learning rate of 5e-5 and optimize the model for 60k iterations with a batch size of 32. The entire training procedure is completed over a period of 40 hours utilizing 8 NVIDIA A100 GPUs.
Figure 4: Qualitative comparisons. We compare the results of several baselines with image prompts and text prompts. In each case, we include the edited portraits as well as their novel view renderings. Our method generates high-quality edited portraits with better 3D quality and alignment with the referenced prompts.
Baselines. We conduct comparisons against two categories of methods that achieve analogous outcomes: 1) 3D reconstruction coupled with 3D editing; 2) 2D editing followed by 3D reconstruction. For the first category, we implement three baseline approaches, wherein we firstly leverage PTI [54] or Live3D [68] for reconstruction and edit both of them following StyleCLIP [49]. Specifically, for experiments in PTI + StyleCLIP, we first conduct pivotal tuning [54] based on the predicted latent codes predicted by e4e [67] to reconstruct the given image in a 3D-aware manner. Then we perform editing on the latent codes following [49], where CLIP loss [53] between the model output and the prompt (no matter it is an image or text) is utilized. The setting of Live3D + StyleCLIP is similar to PTI + StyleCLIP except that we obtain the triplane by Live3D and perform editing on this 3D representation. SPI [79] + StyleGAN-NADA [21] (StyleNADA for abbreviation in Fig. 4) serves as the third baseline in the first category, where SPI is used for 3D reconstruction and StyleGAN-NADA is utilized to optimize the 3D generator for editing. In the second category, we employ a diffusion-based text-guided 2D image editing method IP2P [7] prior to executing 3D reconstruction via Live3D. IP2P being a text-instructed editing model, we need to first extract text prompts from the reference image with the image caption algorithm BLIP [41] when image prompts are input. The 2D edited results then could be obtained by sending the input image and the text prompt to IP2P. The final 3D results are acquired by inverting the 2D editing results with Live3D.
Evaluation Criteria. We evaluate all metrics on 100 pairs of images processed from FFHQ. We conduct a comprehensive evaluation of the editing performance in the following four aspects: identity preservation, reference alignment, 3D consistency, and inference speed. 1) Identity preservation ($\mathbf{ID_{t}}$) aims to measure the preservation of the original identity by calculating the cosine similarity between the identity feature of the input image $\mathbf{I}$ and that of the edited image $\mathbf{I}_{p}$. We use the ArcFace model [16] to extract identity features. 2) Reference alignment ($\mathbf{CLIP_{r}}$) targets at assessing the alignment of the output editing styles to the desired input prompt by computing the cosine similarity in the CLIP [49] feature space. 3) 3D consistency (3D) on the edited outputs is measured following the evaluation protocols established by EG3D [13]. This involves calculating the identity similarity across multiple views. 4) Inference speed (Time) is measured on a single NVIDIA A6000 GPU with an average of 100 samples.
Table 1: Quantitative comparisons. We compare several baselines on the 100 images of FFHQ dataset. It’s important to note that we exclude $CLIP_{r}$ for PTI+StyleCLIP and Live3D+StyleCLIP since these models utilize CLIP for optimization. Our model excels in 3D quality and achieves a remarkable 250x improvement (compared with IP2P+Live3D) in inference speed, achieving real-time performance.
@@ 4.1 Experimental Setup
PTI+StyleCLIP 0.11 - 0.73 360s
Live3D+StyleCLIP 0.63 - 0.72 30s
IP2P+Live3D 0.44 0.62 0.75 10s
@@
Figure 5: The learning process and results of customized prompt adaptation. For the novel style golden statue instructed by reference images, we show the intermediate testing results at 10s, 1min, 2min, and 5min during adaptation with 10 paired training images.
@@
Figure 6: Qualitative comparison for ablations on (a) the distillation loss of 3D GAN ($\mathcal{L}_{3D}$) and (b) diffusion models ($\mathcal{L}_{2D}$).
4.2 Efficient 3D-aware Portrait Editing
We make an in-depth analysis of our efficient portrait editing system both quantitatively and qualitatively. The results are included in Tab. 1 and Fig. 4. For reference alignment, we do not report $CLIP_{r}$ for methods that leverage CLIP for optimization since it is evaluated with CLIP as well. The standout feature of our system is its efficiency. Our method achieves an inference speed of merely 40ms, which improves over 100 times compared to the fastest existing baselines, which require around 10 seconds. Because of the efficient knowledge distillation, our approach is also good in preserving the identity, adhering more closely to the reference prompts, and achieving the best 3D consistency.
Compared with PTI + CLIP methods, our model can enable precise alignment with input and prompts, while the baseline always generates low-fidelity textures, and the resulting geometry has many artifacts. Although the Live3D + CLIP achieves the best $ID_{t}$ score, the edited results are basically unchanged compared to the input. We suspect that Live3D does not have a latent space and loses the editing priors, making it challenging for the results of CLIP optimization to align well with prompts. SPI + StyleGAN-NADA performs relatively slowly (18 minutes per instance) because both the 3D reconstruction and the editing are achieved by optimizing the generator. Although it can produce realistic textures, it fails to reconstruct fine details and to edit following the prompts. In contrast to the 2D editing and subsequent 3D reconstruction pipeline (IP2P + Live3D), our system produces textures consistent with the prompt and aligns better with the input structure. As demonstrated by the instance in the first column, IP2P + Live3D introduces texture artifacts (the secondary face) due to the unstable editing performance of IP2P, while our distillation-based method excels in this aspect. As the second sample in Fig. 4 shows, our model can consistently retain the hair and cloth structure, while IP2P + Live3D cannot handle this. The image caption result by BLIP of the third case turns out to be “a painting of a woman with blue eyes and a red headband”. This sample showcases that, the method of IP2P + Live3D could not handle well editing guided by image prompts due to the improper image caption sent to the 2D editing model, especially for the styles hard to describe with texts. The caption would integrate the content and the desired style of the reference image, which deteriorates both the editing and the identity preservation.
4.3 Adaptation for Customized Editing
@@
With the trained editing network, our system already supports a variety of 3D-aware editing styles. Following DreamBooth [56] and to expand the selection and better conform to users’ preferences, we offer an efficient method for fast adaptation to the user-defined customized prompts. Users can personalize the editing network by providing a modest set of 10 reference editing pairs. These pairs can be either handpicked from artist-created examples or produced using text-guided image editing models. As described in Sec. 3.3, only the partial of the encoder would be optimized. The adaptation process itself is remarkably fast, requiring only about 5 minutes to accomplish the learning of the customized prompts. Please also note that, once the user-defined style has been trained, all the incoming testing results can be obtained in a real-time manner. In Fig. 5, we present the intermediate testing results in the adaptation process. Our model can quickly master the customized knowledge in about 2 minutes. With further training (e.g., 5 minutes), the edited results become more stylized, demonstrating a trade-off between authenticity and stylization. Upon completion of the adaptation, the system allows users to edit testing inputs in these newly learned styles with a minimal inference time of 0.04s. This rapid performance indicates that the system is well-optimized for real-time applications, providing a seamless and efficient user experience.
@@
Table 2: Ablation study on the number of styles used for pretraining. We report LPIPS score after 0.1min, 1min, 2min, and 5min fine-tuning for evaluation.
#Style/Time 0.1min 1min 2min 5min
3 0.7758 0.6370 0.5312 0.5220
10 0.5932 0.5734 0.5106 0.5012
20 0.5703 0.5290 0.5022 0.4849
Table 3: Ablation study on the number of data pairs used for novel prompt adaptation. We report LPIPS score after 0.1min, 1min, 2min, and 5min fine-tuning for evaluation.
@@
Figure 7: Video editing results with 3DPE. We use text prompts "Make him a bronze statue" to edit the input video. Our method can accurately reconstruct and preserve challenging facial expressions and achieve high-quality novel view renderings.
4.4 Ablation Studies
$\mathcal{L}_{2D}$ for Distillation of Diffusion Models. We also study the effect of $\mathcal{L}_{2D}$, which represents using 2D editing priors from the diffusion model [7] to optimize the encoder. From Fig. 6, we observe that $\mathcal{L}_{2D}$ is critical for our model, especially in preserving detailed texture. Without it, the edited image tends to lose the appearance information from prompts and exhibits structural artifacts around the nose.
Ablation on Pretraining Data Scale for Customized Prompt Adaptations. In this part, we conduct ablation studies on the pretraining scale and investigate its influence on customized propmts. Specifically, we choose various 3DPE models pretrained with different style amounts (3, 10, and 20 styles respectively) and perform novel prompt adaptation on a single A6000 GPU. We choose the golden statue as the customized style and report LPIPS score between the predicted stylized images and the pseudo stylized labels produced by IP2P [7]. Experimental results in Tab. 2 reveal that the proposed method could benefit from the improving pretraining data scale and adapt faster and eventually perform better with more data, which suggests the potential of the model for future scaling up.
Ablation on Tuning Data Scale for Customized Prompt Adaptations. To investigate the ability of our method to adapt to user-defined novel prompts, we conduct an ablation study on the number of novel-prompt data pairs required for training. Similar to the prior experiments on the pretraining scale, we perform adaptation with 2, 5, 10, 20, or 50 pairs of novel-prompt data on an A6000 GPU. We experimentally find that our model can converge in about 5 minutes. Therefore, we report the perceptual distance between the model output and the provided ground truth at 0.1min, 1min, 2min, and 5min for comparison. As in Tab. 3, our model can adapt to the new style in around 2 minutes. As the number of provided data pairs increases, the model can adapt faster at the beginning.
4.5 Applications